# Crepant Transformation correspondence For Toric Stack Bundles

### v1

@@ 3.2. Analytic Continuation of $H$-function via Localization
×$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D})},$
@@
×$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}ds.$
@@
[…]
@@
where $D_{j_{-}}\cdot e<0$ and $n$ is a non-negative integer. Choose a contour $C$ such that poles at non-negative integers are on the right of $C$ and other poles are on the left of $C$. In view of [2] and the assumption $\sum D_{j}\in W$, we can see that this integral convergent and analytic as a function of $(y^{+})^{e}$ in the domain $\{(y^{+})^{e}:\lvert\text{arg}((y^{+})^{e})-w\pi\rvert<\pi\}$. Moreover, let $\mathfrak{c}=\prod_{j:D_{j}\cdot e\neq 0}(D_{j}\cdot e)^{D_{j}\cdot e}$ be the so-called conifold point, then for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral is equal to the sum of residues on the right of $C$ and for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$ the integral is equal to negative of the sum of residues on the left of $C$. By our choice of $C$, for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral which equals to the sum of residues on the right of $C$ gives part of the localization of $H^{\mathfrak{D}}_{+}$, and we define the analytic continuation of $H^{\mathfrak{D}}_{+}$ using the sum of residues on the left of $C$ for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$.
To compute the sum of residues on the left, first note that the residues at poles $s=-1-n$ vanish, so we only need to consider poles (3.4). By straight computations, this is
@@
Since $D_{j_{-}}\cdot e<0$, $-D_{j_{-}}\cdot e\in\mathbb{Z}_{+}$. We may write the non-negative integer $n$ as $n=k(-D_{j_{-}}e)+\Lambda_{j_{-}}\mathfrak{D}+l$ with $0\leq l+\Lambda_{j_{-}}\mathfrak{D}<(-D_{j_{-}}e)$. Let $d_{-}=d_{+}+\frac{D_{j_{-}}d_{+}-l}{-D_{j_{-}}e}e$, then
$-\Lambda_{j_{-}}\mathfrak{D}\leq l=D_{j_{-}}\cdot d_{-}<-D_{j_{-}}e-\Lambda_{j_{-}}\mathfrak{D},$
@@
$\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+s(D_{j}\cdot e)=\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)$
@@
Let $f_{-}$ be the equivalent class of $d_{-}$ in $\mathbb{K}_{-}/\mathbb{L}$, then we have relation $d_{+}-d_{-}=f_{+}-f_{-}+Ne$ for some $N\in\mathbb{Z}$. Take this substitution into (3.6) and the dependence on $N$ cancels, so (3.6) is in fact equal to
$-e^{\frac{\sigma_{-}(\delta_{-})}{2\pi i}}((y^{+})^{d_{-}-d_{+}-ke})C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}\prod_{j=1}^{m}\frac{1}{\Gamma(1+\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}.$
Now apply the above result to analytic continuation of $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$. Note the summation over residues on the left of $C$ is the same as a summation over $(\delta_{-},f_{-})$ such that $(\delta_{-},f_{-})|(\delta_{+},f_{+})$. We have
@@
$\mathbb{U}_{H}(\alpha)=\sum_{(\delta,f):\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}}(i^{*}_{(\delta,f)}\alpha)\frac{\mathbf{1}_{\delta,f}}{e_{T}(N_{\delta,f})}+\sum_{(\delta_{+},f_{+}):\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}}\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}(i^{*}_{(\delta_{-},f_{-})}\alpha)\frac{\mathbf{1}_{\delta_{+},f_{+}}}{e_{T}(N_{\delta_{+},f_{+}})}$
@@
[…]
@@
Part (2) of Theorem 0.1 also follows immediately. Recall that $p^{+}_{i}=p^{-}_{i},1\leq i\leq r-1$ form a basis for the wall $W$, so part (2) of Theorem 0.1 is equivalent to $\theta_{+}(p^{+}_{i})\circ\mathbb{U}=\mathbb{U}\circ\theta_{-}(p^{-}_{i})$ for $11\leq i\leq r-$ 1, which follows from definitions of $\theta$ and $\mathbb{U}$.
@@ 3.3. Fourier-Mukai Transformation
Consider the following diagram
${\tilde{\mathcal{P}}}$${\mathcal{P}_{+}}$${\mathcal{P}_{-}}$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$
@@
Recall the fibers $X_{\pm}$ of the toric stack bundle $\mathcal{P}_{\pm}$ are defined from characters $D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}$, stability conditions $\omega_{\pm}\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ and we have the short exact sequence
@@
The toric blow-up $\tilde{X}$ can be defined from $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and a stability condition $\tilde{\omega}$. Consider the following short exact sequence
$0\rightarrow\mathbb{L}\oplus\mathbb{Z}\overset{(\tilde{D}_{1},\dots,\tilde{D}_{m+1})}{\xrightarrow{\hskip 56.9055pt}}\mathbb{Z}^{m+1}\overset{\tilde{\beta}}{\longrightarrow}\mathbf{N}\rightarrow 0$
where $\tilde{D}_{i}$ is defined by
$\tilde{D_{i}}=\begin{cases}D_{i}\oplus 0&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e\leq 0\\ D_{i}\oplus(-D_{i}\cdot e)&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e>0\\ 0\oplus 1&\text{ if }j=m+1\end{cases}$
Let $C_{\pm}$ be the chamber containing $\omega_{\pm}$ and $W$ the wall between chambers. The stability condition $\tilde{\omega}\in(\mathbb{L}\oplus\mathbb{Z})^{\vee}\otimes\mathbb{R}$ is defined as $(\omega_{0},-\epsilon)$ where $\omega_{0}$ is in the relative interior of $W\cap\overline{C_{+}}=W\cap\overline{C_{-}}$ and $\epsilon$ is a small positive real number. By [8, Lemma 6.16], the toric stack $\tilde{X}$ defined from the above data is the common toric blow-up of $X_{\pm}$.
$L(p)=(P\times_{(\mathbb{C}^{\times})^{m}}U_{\omega})\times\mathbb{C}/(b,z,v)\sim(b,gz,g(p)v)$
where $g\in K,b\in B,z\in U_{\omega},v\in\mathbb{C}$. This line bundle defines a class in $K^{0}_{T}(\mathcal{P})$. Especially, let $p=D_{i}$, we define the following class in $K^{0}_{T}(\mathcal{P}_{\pm})$
$R^{\pm}_{i}=L_{\pm}(D_{i})\otimes e^{\lambda_{i}}$
Similarly we may define line bundle $L(p,n)$ over $\tilde{\mathcal{P}}$ for a character $(p,n)\in\text{Hom}(K\times\mathbb{C}^{\times},\mathbb{C}^{\times})=\mathbb{L}^{\vee}\oplus\mathbb{Z}$, and classes $\tilde{R}_{i}=L(\tilde{D}_{i})\otimes e^{\lambda_{i}}$ for $1\leq i\leq m$ and $\tilde{R}_{m+1}=L(\tilde{D}_{m+1})=L(0,1)$ in $K^{0}_{T}(\tilde{\mathcal{P}})$. Note we have $f^{*}_{-}(L_{-}(p))=L(p,0)$ and $f^{*}_{+}(L_{+}(p))=L(p,-p\cdot e)$. If we let $k_{i}=\text{max}(D_{i}\cdot e,0)$ and $l_{i}=\text{max}(-D_{i}\cdot e,0)$, especially we have $f^{*}_{-}R^{-}_{i}=\tilde{R}_{i}\tilde{R}_{m+1}^{k_{i}}$ and $f^{*}_{+}R^{+}_{i}=\tilde{R}_{i}\tilde{R}_{m+1}^{l_{i}}$.
@@
$\left\{e_{\delta,\varrho}:=(i_{\delta})_{*}\varrho=L(\hat{\varrho})\prod_{i\notin\delta}(1-S_{i})\right\}$
@@
Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{\tilde{\omega}}$ for $\tilde{\mathcal{P}}$ with all $j_{1},\dots,j_{r-1},j_{-},j_{+}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Then we have the following diagram
[…]
@@
$G_{\tilde{\delta}}$ =$=\{(g,\lambda)\in K\times\mathbb{C}^{\times}:D_{j_{-}}(g)=1,D_{j}(g)\lambda^{-D_{j}\cdot e}=1\ \text{for}\ j\in\delta_{+}\}$
$G_{\delta_{+}}$ =$=\{h\in K:D_{j}(h)=1\ \text{for all}\ j\in\delta_{+}\}.$
@@
$1\rightarrow\boldsymbol{\mu}_{l}\rightarrow G_{\tilde{\delta}}\rightarrow G_{\delta_{+}}\rightarrow 1$
@@
where $\tilde{P}$ is a principal (×$(\mathbb{C}^{\times})^{m+1}$-bundle, $U_{\tilde{\omega}}$ is defined by the GIT data $\tilde{\omega}$, $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and $\mathcal{A}_{\tilde{\omega}}$ as introduced in Section 2.2, and for any $(g,\lambda)\in K\times\mathbb{C}^{\times}$, $z=(z_{1},\dots,z_{m+1})\in U_{\tilde{\omega}}$, the action is defined by $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$, especially we have
@@
$\frac{1}{l}\sum_{t\in\mathcal{T}}t^{n}=\left\{\begin{matrix}(R^{+}_{j_{-}})^{n/l}\ &\text{if}\ l\ \text{divides}\ n;\\ 0&\text{otherwise}\end{matrix}\right.$
To prove the lemma, it suffices to consider monomials $q(\tilde{R}_{1},\dots,\tilde{R}_{m+1})=\prod_{i=1}^{m+1}\tilde{R}_{i}^{r_{i}}$:
@@
$\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho}$
where $\delta$ is regarded as a minimal anticone in $\mathcal{A}_{-}$ on the left and in $\mathcal{A}_{+}$ on the right. Otherwise for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$ we have
$\mathbb{FM}(e_{\delta_{-},\varrho})=\frac{1}{l}\sum_{t\in\mathcal{T}}\frac{1-S^{+}_{j_{-}}}{1-t^{-1}}L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\prod_{i\notin\delta_{-}}(1-t^{-k_{i}}S^{+}_{i})$
@@
For $\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}$, the birational map $\varphi:\mathcal{P}_{+}\dashrightarrow\mathcal{P}_{-}$ is an isomorphism in a neighborhood of $B_{\delta}$, so $\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho}$. Now for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$, let $\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}$ be such that $\delta_{+}|\delta_{-}$. Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{~{}}$ for $\tilde{\mathcal{P}}$ with all $\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Recall the following diagram
[…]
@@
Apply Lemma 3.4, we have
$(f_{+})_{*}(f_{-})^{*}e_{\delta_{-},\varrho}$ $=\sum_{\delta_{+}:\delta_{+}|\delta_{-}}(i_{\delta_{+}})_{*}(i_{\delta_{+}})^{*}(\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\frac{\prod_{i\notin\delta_{-}}(1-t^{l_{i}-k_{i}}S^{+}_{i})}{(1-t^{-1})\prod_{i\notin\tilde{\delta}}(1-t^{l_{i}}S^{+}_{i})})$
@@
$\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\frac{1-S^{+}_{j_{-}}}{1-t^{-1}}\prod_{i\notin\delta_{-}}(1-t^{-k_{i}}S^{+}_{i})$
@@
First recall that $(i_{\delta})^{*}S^{+}_{i}=1$ for $i\in\delta$, so for a minimal anticone $\delta$, if there is $i\in\delta$ but $i\notin\delta_{-}$ such that $D_{i}\cdot e\leq 0$, then the restriction of $\prod_{i\notin\delta_{-}}(1-t^{-k_{i}}S^{+}_{i})$ vanishes so the restriction of above summation is 00. Another case is that $j_{-}\in\delta$ and there is some $j_{+}\in\delta$ such that $D_{j_{+}}\cdot e>0$. Then for each $t\in\mathcal{T}$, there is a factor

### v2

@@ 3.2. Analytic Continuation of $H$-function via Localization
$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D})},$
@@
$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}ds.$
@@
$e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}\int_{C}\Gamma(s)\Gamma(1-s)\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}(e^{-\pi i\omega}(y^{+})^{e})^{s}ds.$
@@
where $D_{j_{-}}\cdot e<0$ and $n$ is a non-negative integer. Choose a contour $C$ such that poles at non-negative integers are on the right of $C$ and other poles are on the left of $C$. In view of [2] and the assumption $\sum_{j}D_{j}\in W$, we can see that this integral is convergent and analytic as a function of $(y^{+})^{e}$ in the domain $\{(y^{+})^{e}:\lvert\text{arg}((y^{+})^{e})-w\pi\rvert<\pi\}$. Moreover, let $\mathfrak{c}=\prod_{j:D_{j}\cdot e\neq 0}(D_{j}\cdot e)^{D_{j}\cdot e}$ be the so-called conifold point, then for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral is equal to the sum of residues on the right of $C$ and for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$ the integral is equal to the negative of the sum of residues on the left of $C$. By our choice of $C$, for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral, which is equal to the sum of residues on the right of $C$, gives part of the localization of $H^{\mathfrak{D}}_{+}$, and we define the analytic continuation of $H^{\mathfrak{D}}_{+}$ using the sum of residues on the left of $C$ for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$.
To compute the sum of residues on the left, first note that the residues at poles s=1ns=-1-n vanish, so we only need to consider poles (3.4). By straightforward computations, this is
@@
Since $D_{j_{-}}\cdot e<0$, $D_{j_{-}}\cdot e\in\mathbb{Z}_{+}$, we may write the non-negative integer $n$ as $n=k(-D_{j_{-}}e)+\Lambda_{j_{-}}\mathfrak{D}+l$ with $0\leq l+\Lambda_{j_{-}}\mathfrak{D}<(-D_{j_{-}}e)$. Let $d_{-}=d_{+}+\frac{D_{j_{-}}d_{+}-l}{-D_{j_{-}}e}e$, then
$\Lambda_{j_{-}}\mathfrak{D}\leq l=D_{j_{-}}\cdot d_{-}<-D_{j_{-}}e-\Lambda_{j_{-}}\mathfrak{D},$
@@
$\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+s(D_{j}\cdot e)=\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e),$
@@
Let $f_{-}$ be the equivalent class of $d_{-}$ in $\mathbb{K}_{-}/\mathbb{L}$, then we have the relation $d_{+}-d_{-}=f_{+}-f_{-}+Ne$ for some $N\in\mathbb{Z}$. Take this substitution into (3.6) and the dependence on $N$ cancels, so (3.6) is in fact equal to
$e^{\frac{\sigma_{-}(\delta_{-})}{2\pi i}}((y^{+})^{d_{-}-d_{+}-ke})C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}\prod_{j=1}^{m}\frac{1}{\Gamma(1+\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}.$
Now apply the above result to analytic continuation of $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$. Note that the summation over residues on the left of $C$ is the same as the summation over $(\delta_{-},f_{-})$ such that $(\delta_{-},f_{-})|(\delta_{+},f_{+})$. We have
@@
$\mathbb{U}_{H}(\alpha)=\sum_{(\delta,f):\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}}(i^{*}_{(\delta,f)}\alpha)\frac{\mathbf{1}_{\delta,f}}{e_{T}(N_{\delta,f})}+\sum_{(\delta_{+},f_{+}):\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}}\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}(i^{*}_{(\delta_{-},f_{-})}\alpha)\frac{\mathbf{1}_{\delta_{+},f_{+}}}{e_{T}(N_{\delta_{+},f_{+}})}.$
@@
(3.7) $H^{\bullet\bullet}_{T}(I\mathcal{P}_{-})\otimes_{\widehat{R}_{T}}\widehat{S}_{T}$$H^{\bullet\bullet}_{T}(I\mathcal{P}_{+})\otimes_{\widehat{R}_{T}}\widehat{S}_{T}$$\tilde{\mathcal{H}}(\mathcal{P}_{-}):=H^{\bullet}_{CR,T}(\mathcal{P}_{-})\otimes_{R_{T}}S_{T}[\text{log}z]((z^{-1/k}))$$H^{\bullet}_{CR,T}(\mathcal{P}_{+})\otimes_{R_{T}}S_{T}[\text{log}z]((z^{-1/k}))=:\tilde{\mathcal{H}}(\mathcal{P}_{+}).$$\scriptstyle{\mathbb{U}_{H}}$$\scriptstyle{\tilde{\Psi}_{-}}$$\scriptstyle{\tilde{\Psi}_{+}}$$\scriptstyle{\mathbb{U}}$
@@
Part (2) of Theorem 0.1 also follows immediately. Recall that $p^{+}_{i}=p^{-}_{i},1\leq i\leq r-1$ form a basis for the wall $W$, so part (2) of Theorem 0.1 is equivalent to $\theta_{+}(p^{+}_{i})\circ\mathbb{U}=\mathbb{U}\circ\theta_{-}(p^{-}_{i})$ for $1\leq i\leq r-1$, which follows from definitions of $\theta$ and $\mathbb{U}$.
@@ 3.3. Fourier-Mukai Transformation
Consider the following diagram:
$\tilde{\mathcal{P}}$$\mathcal{P}_{+}$$\mathcal{P}_{-},$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$
@@
Recall that the fibers $X_{\pm}$ of the toric stack bundle $\mathcal{P}_{\pm}$ are defined from characters $D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}$, stability conditions $\omega_{\pm}\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ and we have the short exact sequence:
$0\rightarrow\mathbb{L}\overset{(D_{1},\dots,D_{m})}{\xrightarrow{\hskip 28.45274pt}}\mathbb{Z}^{m}\overset{\beta}{\longrightarrow}\mathbf{N}\rightarrow 0.$
The toric blow-up $\tilde{X}$ can be defined from $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and a stability condition $\tilde{\omega}$. Consider the following short exact sequence:
$0\rightarrow\mathbb{L}\oplus\mathbb{Z}\overset{(\tilde{D}_{1},\dots,\tilde{D}_{m+1})}{\xrightarrow{\hskip 56.9055pt}}\mathbb{Z}^{m+1}\overset{\tilde{\beta}}{\longrightarrow}\mathbf{N}\rightarrow 0,$
where $\tilde{D}_{i}$ is defined by
$\tilde{D_{i}}=\begin{cases}D_{i}\oplus 0&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e\leq 0,\\ D_{i}\oplus(-D_{i}\cdot e)&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e>0,\\ 0\oplus 1&\text{ if }j=m+1.\end{cases}$
Let $C_{\pm}$ be the chamber containing $\omega_{\pm}$ and $W$ the wall between the two chambers. The stability condition $\tilde{\omega}\in(\mathbb{L}\oplus\mathbb{Z})^{\vee}\otimes\mathbb{R}$ is defined as $(\omega_{0},-\epsilon)$ where $\omega_{0}$ is in the relative interior of $W\cap\overline{C_{+}}=W\cap\overline{C_{-}}$ and $\epsilon$ is a small positive real number. By [8, Lemma 6.16], the toric stack $\tilde{X}$ defined from the above data is the common toric blow-up of $X_{\pm}$.
$L(p)=(P\times_{(\mathbb{C}^{\times})^{m}}U_{\omega})\times\mathbb{C}/(b,z,v)\sim(b,gz,g(p)v),$
where $g\in K,b\in B,z\in U_{\omega},v\in\mathbb{C}$. This line bundle defines a class in $K^{0}_{T}(\mathcal{P})$. Especially, let $p=D_{i}$, we define the following class in $K^{0}_{T}(\mathcal{P}_{\pm})$:
$R^{\pm}_{i}=L_{\pm}(D_{i})\otimes e^{\lambda_{i}},$
@@
$\left\{e_{\delta,\varrho}:=(i_{\delta})_{*}\varrho=L(\hat{\varrho})\prod_{i\notin\delta}(1-S_{i})\right\}.$
@@
Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{\tilde{\omega}}$ for $\tilde{\mathcal{P}}$ with all $j_{1},\dots,j_{r-1},j_{-},j_{+}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Then we have the following diagram:
$B_{\delta_{-}}$$\mathcal{P}_{-}$$B_{\tilde{\delta}}$$\tilde{\mathcal{P}}$$B_{\delta_{+}}$$\mathcal{P}_{+}.$$\scriptstyle{i_{\delta_{-}}}$$\scriptstyle{f_{-,\tilde{\delta}}}$$\scriptstyle{f_{+,\tilde{\delta}}}$$\scriptstyle{i_{\tilde{\delta}}}$$\scriptstyle{f_{-}}$$\scriptstyle{f_{+}}$$\scriptstyle{i_{\delta_{+}}}$
@@
$G_{\tilde{\delta}}$ $=\{(g,\lambda)\in K\times\mathbb{C}^{\times}:D_{j_{-}}(g)=1,D_{j}(g)\lambda^{-D_{j}\cdot e}=1\ \text{for}\ j\in\delta_{+}\},$
$G_{\delta_{+}}$ $=\{h\in K:D_{j}(h)=1\ \text{for all}\ j\in\delta_{+}\}.$
@@
$1\rightarrow\boldsymbol{\mu}_{l}\rightarrow G_{\tilde{\delta}}\rightarrow G_{\delta_{+}}\rightarrow 1.$
@@
where $\tilde{P}$ is a principal $(\mathbb{C}^{\times})^{m+1}$-bundle, $U_{\tilde{\omega}}$ is defined by the GIT data $\tilde{\omega}$, $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and $\mathcal{A}_{\tilde{\omega}}$ as introduced in Section 2.2, and for any $(g,\lambda)\in K\times\mathbb{C}^{\times}$, $z=(z_{1},\dots,z_{m+1})\in U_{\tilde{\omega}}$, the action is defined by $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$, especially we have
@@
$\frac{1}{l}\sum_{t\in\mathcal{T}}t^{n}=\left\{\begin{matrix}(R^{+}_{j_{-}})^{n/l}\ &\text{if}\ l\ \text{divides}\ n,\\ 0&\text{otherwise}.\end{matrix}\right$.
To prove the lemma, it suffices to consider the case of monomials $q(\tilde{R}_{1},\dots,\tilde{R}_{m+1})=\prod_{i=1}^{m+1}\tilde{R}_{i}^{r_{i}}$:
@@
$\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho},$
where $\delta$ is regarded as a minimal anticone in $\mathcal{A}_{-}$ on the left and in $\mathcal{A}_{+}$ on the right. Otherwise, for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$ we have
[…]
@@
For $\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}$, the birational map $\varphi:\mathcal{P}_{+}\dashrightarrow\mathcal{P}_{-}$ is an isomorphism in a neighborhood of $B_{\delta}$, so $\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho}$. Now for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$, let $\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}$ be such that $\delta_{+}|\delta_{-}$. Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{~}$ for $\tilde{\mathcal{P}}$ with all $\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Recall the following diagram:
$B_{\delta_{-}}$$\mathcal{P}_{-}$$B_{\tilde{\delta}}$$\tilde{\mathcal{P}}$$B_{\delta_{+}}$$\mathcal{P}_{+}.$$\scriptstyle{i_{\delta_{-}}}$$\scriptstyle{f_{-,\tilde{\delta}}}$$\scriptstyle{f_{+,\tilde{\delta}}}$$\scriptstyle{i_{\tilde{\delta}}}$$\scriptstyle{f_{-}}$$\scriptstyle{f_{+}}$$\scriptstyle{i_{\delta_{+}}}$
@@
Applying Lemma 3.4, we have
[…]
@@
[…]
@@
[…]